OpenStudy (anonymous):

If |8i-j|=|7i+kj| (vectors), find two possible values for k belongs to R. Answers: +4, -4. How?

OpenStudy (amistre64):

becasue it is asking for the length of the vectors to be equal

OpenStudy (anonymous):

@amistre64 I need step by step how to do it... I undestand why it is +4 and -4

OpenStudy (amistre64):

\[8^2+1^2=7^2+k^2\] \[8^2+1^2-7^2=k^2\]

OpenStudy (anonymous):

Oooooooooh my God! It's that simple! Thank you very much!

OpenStudy (amistre64):

youre welcome